Find the particular integral of dy/dx+y=cos3x
1
Expert's answer
2019-05-15T14:19:53-0400
"y'+y= cos(3x)"


Consider:

"y(x)=u(x)*v(x)\\
u'v+v'u+uv=cos(3x)\\
uv'+v(u'+u)=cos(3x)\\"

then:


"\frac{du}{dx}=-u\\
\frac{du}{u}=-dx\\
\int \frac{du}{u}=-\int dx\\
ln(u)=-x \\
u=e^{-x}"

then:

"v'u=cos(3x)\\
v'=cos(3x)*e^{x}\\
v=\int cos(3x)*e^x=\frac{1}{10}cos(3x)*e^x+\frac{3}{10}sin(3x)*e^x+C\\"

then a particular integral is


"y(x)=\frac{1}{10}cos(3x)+\frac{3}{10}sin(3x)"